how to find sides and angles of pedal triangle??

Grade:12th pass

1 Answers

Apoorva Arora IIT Roorkee
9 years ago
Given a point P inside the triangle ABC, The pedal triangle of P is the triangle whose polygon vertices are the feet of the perpendiculars from P to the sidelines. The pedal triangle of a triangle with trilinear coordinates\alpha :\beta :\gammaand angles A B and C has trilinear vertex matrix
\begin{bmatrix} 0 &\beta +\alpha cosC &\gamma +\alpha cosB \\ \alpha +\beta cosC& 0 &\gamma +\beta cosA \\ \alpha +\gamma cosB & \beta +\gamma cosA & 0 \end{bmatrix}
The side lengths are,
a'=\frac{abc\sqrt{\beta ^{2}+\gamma ^{2}+2\beta \gamma cosA}}{2R|a\alpha +b\beta +c\gamma |}
b'=\frac{abc\sqrt{\alpha ^{2}+\gamma ^{2}+2\alpha \gamma cosB}}{2R|a\alpha +b\beta +c\gamma |}
c'=\frac{abc\sqrt{\alpha ^{2}+\beta ^{2}+2\alpha \beta cosC}}{2R|a\alpha +b\beta +c\gamma |}
where R is the circumradius of the triangle ABC.
Thanks and Regards
Apoorva Arora
IIT Roorkee
askIITias Faculty